OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are 40 and 5000, respectively.

OpenStudy (anonymous):

@satellite73 can you help with this?

OpenStudy (anonymous):

@Apprentice I'm afraid your face is on the keyboard and you're sleeping lol

OpenStudy (anonymous):

Here are the answer choices: an = 8 • 5n - 2 an = 8 • 5n + 3 an = 8 • 5n - 1 an = 8 • 5n + 1

OpenStudy (anonymous):

is it the 3rd one?